What happens in the relativistic addition of velocities formula (derived in
problem (10) when one of the velocities (say, vBA) is the speed of light??
1
Expert's answer
2019-11-11T15:56:49-0500

relativistic addition of velocities formula:


"v_B = \frac{v_A+v_{BA}}{1+v_A \cdot v_{BA}}"

If one of the velocities (say, vBA) is the speed of light, "v_{BA}=1" and

"v_B = \frac{v_A+1}{1+v_A \cdot 1} = 1"

Therefore, "v_B = 1" and does not depend on "v_A" .